OpenStudy (unklerhaukus):

\[\delta_{ij}\sum_{i=12}^{12}\prod_{k=j}^{k}j^2\]

OpenStudy (anonymous):

what do you mean by k=j to k ?

OpenStudy (unklerhaukus):

j=k*

OpenStudy (anonymous):

and i = 12 to 12 ? :O

OpenStudy (unklerhaukus):

thats right

OpenStudy (anonymous):

so as I see it: there is not really summing nor multiplying .. it looks like k = 12 and that is it. so 12^2 = 144 (i = j otherwise the delta is zero)

OpenStudy (unklerhaukus):

you got it !

OpenStudy (anonymous):

:)

hartnn (hartnn):

\(\sum \limits_{i=12}^{12}k^2=k^2\) right ?

hartnn (hartnn):

so, \(\delta_{ij}k^2=0\) so, 0 is final answer ?? just confirming...

OpenStudy (unklerhaukus):

notice i made an error when i wrote out the question the first time i put k=j when i meant j=k

hartnn (hartnn):

\(\delta_{ij}\sum_{i=12}^{12}\prod_{j=k}^{k}j^2=\delta_{ij}\sum_{i=12}^{12}k^2=\delta_{ij}k^2=0\) ?

OpenStudy (unklerhaukus):

hmm i guess i should have left it as k=j

hartnn (hartnn):

no the variable is j, so it must be j= something to something

hartnn (hartnn):

i don't get how its 144

hartnn (hartnn):

@Coolsector , how is k=12 ??

OpenStudy (unklerhaukus):

the variable doesn't have to appear in the argument